A company that makes handheld computers is trying to improve one of its poorer selling products. They conducted some customer surveys and found that customers prefer the products to weigh between 10 and 14 ounces. At present, the mean of the process is 12 ounces and the standard deviation of the process is .4 ounces. They want to ensure that the production process will meet customer expectations; that is, the process capability (Cp) will be greater than or equal to 1.00. a. What is the Cp of the current production process? b. What does the standard deviation of the process have to be for the Cp to equal 1.00?
Answer #1

a. Cp= (UTL - LTL)/[(6)(St dev of process)] = (14 - 10)/[(6)(.4)] = 4/2.4 = 1.67  

Current process meets customer expectations.    

b. Std. dev. of process = (UTL - LTL)/[(6)(Cp)] = (14 - 10)/[(6) (1. 0)] = 0.667